Solve the following word problems. Be sure to show the equation you use for the solution
56) Science and medicine. A bus leaves a station at 1 P.M., traveling west at an average rate of 44 mi/h. One hour later a second bus leaves the same station, traveling east at a rate of 48 mi/h. At what time will the two buses by 274 mi apart?

Answer by stanbon(48568)

A bus leaves a station at 1 P.M., traveling west at an average rate of 44 mi/h. One hour later a second bus leaves the same station, traveling east at a rate of 48 mi/h. At what time will the two buses by 274 mi apart?
---------------
1st bus DATA;
rate = 44 mph ; time = x hrs ; distance = 44x miles
-------------------
2nd bus DATA:
rate = 48 mph ; time = x-1 hrs ; distance = 48x-48 miles
--------------------
EQUATION:
distance + distance = 274
44x + 48x -48 = 274
92x = 322
x = 3.5 hrs
1 P.M. + 3.5 hrs = 4:30 P.M.
===============
Cheers,
Stan H.